OpenStudy (anonymous):

A furniture maker uses the specification 21.88 < w

OpenStudy (anonymous):

\[\left| w-22 \right|<0.12\]

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

|w−δ|<ϵ where ϵ=Tolerence The solution to this enequality is ( similar form given in your problem) −ϵ+δ<w<ϵ+δ to find δ add −ϵ+δ and ϵ+δ and then divide by 2 = 22 to find ϵ subtract ϵ+δ and −ϵ+δ and then divide by 2 = 0.12 then |w−δ|<ϵ |w−22|<0.12